GreenHab Report – April 30th

GreenHab Report

Date: 30Apr2018

Environmental control: Ambient with door open

Shade cloth: on

Fan: not used today

Average temperatures: 24.3 ℃

Low: 15.5℃

High: 50.2℃

Hours of supplemental light: None

Changes to crops: Yellowing resolving, plants seem to be in slightly better health.

Daily water usage for crops: 7.25 Gallons

Time(s) of watering for crops: Twice: In the morning at 0800 and at 1845.

Morning research observations: None

Changes to research plants: None

Aquaponics: None

Narrative: Yellowed plants are are turning green again after extra watering. It may or may not help that we are talking to them gently but lovingly.

Support/supplies needed: None

Operations Report – April 30th

Crew 193 Operations Report 30Apr2018

SOL: 8

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10. Soldering Irons

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now). At 4:45 PM today, the SoC was 73%, which is lower than it should be, given how sunny it is today; there are some storm clouds to the southwest, which could be blocking the sun, we need to keep an eye on SoC.
I will repair charging plugs once we have a soldering iron.

Generator (hours run): 10:45 pm to 8:10 am.

Solar— SOC 59 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0.5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134370.9

Toilet tank emptied: no
Date last emptied: April 29

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 300, 350.1, 350.3

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 2 Gallons

# Hours the ATVs were Used today: 3

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.

Summary of suits and radios: Still work to do on two piece suits. Repair of suit chargers still in progress. Other than #10, one piece suits nominal.

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment proceeding.

Summary of RAM operations: A charging connector was moved from RAM for space suit repair (once soldering iron arrives)

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: We would like some black straps used for connecting two piece suit neck rings to backpacks. I did not see any in the RAM.

Science Report – April 30th

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no new sites were noted as useful for emergency evacuation shelters. The second EVA found a location in Candor Chasma that had been previously scouted by the other EVA team last week.

2. Sunspot and Solar Flare Monitoring: The crew astronomer saw one filament on the sun. No other features were apparent.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: Both EVAs took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water.

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection. No fogging was ever reported.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Commander’s Report – April 30th

Commander, PHEnOM Gold Crew, MDRS #193

Mission Sol 8 (Earth Day April 30th, 2018)

Exactly a week ago we arrived at Mars and embarked on this mission, excited, curious, and eager to learn, explore, research and contribute to mankind’s endeavor to colonize Mars. It has been a rewarding week what with all our research progressing well, exploration and scouting for emergency shelters giving positive results and barring a few minor incidents, the crew staying fit and fine.

The crew has adapted well to the harsh Martian conditions and is functioning great as one unit. I am impressed at how each member is being creative and innovative in coming up with solutions to problems and willing to take on new projects to help in every way to make this mission a success.

Mars is challenging ; its comforting to see the crew maintain a stable emotional quotient and in great psychological health even after a week here! Spirits remain high and the crew looks forward to completing their mission goals and more.

We have been working hard and playing hard. Games, movies, a lot of sharing stories and learning ASL have been our fun pastimes. Our crew is a great example of how people from different backgrounds, age groups and professions can come together and work as a team to make a mission successful.

Every sol on Mars has been an adventure to behold,
For this outstanding crew, call us 193 or Gold!
With all the work we do here, we’ll leave our mark on Mars,
As we continue to further mans mission, to travel to the stars !
– Ur Poet Commander 😉

Journalist Report – April 29th

Journalist Report – Earth-Date 29APR2018 / Mars Sol 7

Author: Eric Shear

This weekend, we took a much-needed rest day on Sol 6. That doesn’t mean we lounged around all day – we made outreach videos with a combination of spoken and signed language.

Today (Sol 7) began the first of our switched EVAs. Team Canada had previously done morning EVAs, and is now doing EVAs in the afternoon, and vice versa for the Americans.

On Sol 6, our geologist/astronomer finally caught a prominence on the Sun, much to the excitement of the astronomers at Mission Support. Then he did it again on Sol 7, getting three solar features despite heavy cloud cover. We are glad to see him sharpen his already considerable telescope and image processing skills.

The morning of Sol 7, Team America went on EVA to Zubrin’s Head, White Rock Canyon, and Kissing Camel. Team Canada explored Toothy Ridge and Copernicus Highway. During one of these EVAs, a crewmember injured her knee by falling down a hill.

The afternoon was spent recuperating, working on projects and furthering our research. A quiet evening of eating good food and low-key recreation is anticipated. Our XO is anxious for a rematch in cards…

Astronomy Report – April 29th

Astronomy Report
Name: Eric Shear
Crew: 193
Date: April 29, 2018

Sky Conditions: Cloudy with sun still intense through them

Wind Conditions: Windy around noon/afternoon

Observation Start Time: 1315

Observation End Time: 1415

Summary:

Two dark features and a white feature were spotted on the sun today. The dark features are not defined enough to look like sunspots, so they could be filaments. The white spot is almost definitely another solar flare, smaller than the one spotted two sols ago.

Photos were taken with the Skyris camera and the best images processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The results are attached.

Objects Viewed: Sun

Problems Encountered:

Winds were such that I could not open the lower shutter without it shaking too much, so I kept it closed. The sun was high enough in the sky by then that I could observe it with just the upper shutter retracted.

Passing clouds blotted out the sun at times, so I waited until the camera image cleared before taking captures.

EVA Report – April 29th

Sol 7 Crew 193 EVA 13 & 14 Summary
Earth-Date 29APR2018

EVA 13

EVA Gold Crew Members – Gold000 (CO Patil-Sabale) / Gold 003 (CE Attig) / Gold005 (CGA Shear)

EVA Objectives – Summary
1. Explored White Rock Canyon, and Kissing Camel Ridge. Found a few potential shelter sites, and a small pond of muddy water!
2. Repeated GPS recording experiment.

3. Continue to explore the crew biometrics study while on EVA. Raw data here
4. Continuing visor fogging study. Raw data here

Issues/Mitigation

  • Was super windy. No major issues

ATVs/Rovers Used

  • ATV 300
  • Honda ATV
  • ATV #350.3

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
  • Adequate hydration – No
  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

EVA 14

EVA Gold Crew Members – Gold001 (XO Campbell) / Gold 004 (HSO Pandya) /

EVA Objectives – Summary

  1. Explore Toothy Ridge Area and Copernicus hwy
  2. Continue to explore the crew biometrics study while on EVA / Raw data here
  3. Continued search for emergency shelters
  4. Continuing visor fogging study / Raw data here

Issues/Mitigation

  • Extreme winds.

ATVs/Rovers Used
Honda ATV
ATV #350.3

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
  • Adequate hydration – No
  • Adequate protection – Flightsuit/footwear/visor/glov e precautions – Yes

Warm Regards!
Commander, (Gold 00), Crew #193

Mars Desert Research Station

Science Report – April 29th

Science Report for 29 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no sites were noted as useful for emergency evacuation shelters. If a crew was trapped north on Copernicus highway, there are several low sites where they could shelter out of the wind.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found two dark features on the sun and another possible solar flare.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: The first EVA of the day took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Final samples of the project were taken from a towel wiped dirty dish that had been place under UV and a dirty dish that had been wiped with dirt and then placed under a UV light. Bacteria growth and monitoring will continue over the remaining days of the mission.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: No fogging was reported.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Thanks,
Doug
Crew 193 XO and Science Officer